# Muharaf: Manuscripts of Handwritten Arabic Dataset for Cursive Text Recognition

### v1

@@ 1 Introduction
The rest of this paper is organized as follows. Section 2 overviews the existing Arabic datasets for OCR/HTR. Section 3 describes the pipeline for collecting data and Section 4 details the dataset features, formats, and statistics. Section 5 provides the baseline results from our preliminary experiments on HTR. Section 6 concludes the paper and discusses the limitations of this work along with future directions.
@@ 2 Publicly available Arabic resources
The number of publicly available Arabic datasets is fewer than those available for languages written in the Latin script. Many Arabic datasets were tailored for specialized tasks, e.g., BADAM for baseline detection [11], HADARA80P for word spotting [22], AHDB for detecting and recognizing numbers on legal checks [2], and WAHD for writer identification [1]. In our subsequent discussion, we focus on offline, handwritten datasets for Arabic HTR, where ground truths for text are available. We may classify offline handwritten HTR datasets into two categories:
@@
[separator=pipe, tabular=C1.7cm C1.3cm C1.3cm L7.7cm, table head=Dataset Category Annotated Text Lines Dataset Description
@@
Table 2: A comparison of various HTR datasets.
@@ 3 Data collection process
We developed an image-labeling software named ScribeArabic specialized for annotating and transcribing Arabic page images of the Muharaf dataset. This software allows a user to annotate text lines in a browser window and transcribe them in a panel next to it. A separate module has the option to label various page elements. A screenshot of this software is shown in Figure 2. We plan to make both the source code for ScribeArabic and the Muharaf dataset publicly available.
@@
We assembled a team of expert Arabic speakers to transcribe historic Arabic manuscripts with technical support from machine learning researchers. The text lines of the first 180 images of our curated dataset were annotated by non-Arabic speakers using the ScribeArabic software. A Lebanese Arabic history professor then manually entered the transcriptions into an Excel sheet. We chose Excel for entering transcriptions because initially ScribeArabic supported only line annotations. For the next 1400140014001400+ images, we used our upgraded ScribeArabic software to allow the direct input of transcriptions in a browser window. The annotations and transcriptions for these 1400140014001400+ images were performed by two native Lebanese Arabic speakers adept at archival Arabic. Their transcriptions were checked by a third Lebanese Arabic expert who was also a historian.
@@ 4.1 PAGE-XML format for OCR/HTR datasets
The PAGE-XML format is an XML-based page image representation framework introduced by the PRImA Research Lab [26]. This format incorporates various image characteristics as well as the information on page layout and its contents at various levels of granularity. The official, full description of the XML schema is hosted at https://www.primaresearch.org. A conceptual class diagram of the subset of the PAGE-XML hierarchy used in the Muharaf dataset is shown in Figure 3. Where applicable, the following labeled regions are present in the PAGE-XML file of each image:
@@ 5 Baseline HTR results and discussion
Table 3: HTR results repeated over three random splits of Muharaf
@@ 5.1 Evolution of HTR performance as more data were collected
6 Conclusions and future directions
@@
To the best of our knowledge, Muharaf is the largest publicly available Arabic dataset comprising fully annotated and transcribed historical manuscript pages at the text-line level. However, this collection is not devoid of limitations. The process of identifying all the writers and the exact timeline for each document can continue after the initial release of the dataset. For documents where the writer’s information cannot be extracted, e.g., the case where a scribe penned a legal document or a church record, one may want to define categories of different writers and writing styles.
Another area of interest is the use of Muharaf transcriptions for the extraction of linguistic knowledge and the identification of the colloquial form of the Arabic language used in a particular period. Language models based on this information may improve the performance of the HTR system. We invite researchers in the document analysis and OCR community to utilize this dataset and advance the state of the art.
@@
We thank Stephen Randall Filios from Family Search for initiating discussions and providing feedback on tagging page elements in document images. We thank Elham Abdallah for providing support and coordinating work between NC State and USEK.
@@ Appendix C Procedure for generating keywords for each collection
You will be given Arabic text
Based on the text, assign it the most three relevant keywords from this list:
@@ 4. Give output as JSON format with keys:
Note that the prompt and the keywords present above were manually refined iteratively based on the output of GPT.
@@ D.3 PAGE-XML converter and page elements viewer
We include the code for converting ScribeArabic’s JSON files to PAGE-XML files using a PAGE-XML converter. Instructions for downloading the code and running it are at https://github.com/mehreenmehreen/xml_converter. Users can also download a custom viewer for inspecting the annotated page elements on a document page. The viewer shows the different annotated page elements in different colors. A screenshot of this viewer is shown in Figure 6. This app was written in Python using the Tkinter library.
Figure 6: Viewer of the annotated page elements showing floating regions, a paragraph region, and a signature region. Individual text lines are enclosed in black polygonal boundaries.
@@ D.4 Start, Follow, Read — Arabic
  • Additional inference results from training the Muharaf dataset and its public portion.
D.5 License
@@
Appendix E Datasheet for Muharaf Dataset
@@ E.1 Motivation
The dataset was created from a collaboration between:
  • A team of historians and archivists at the Phoenix Center for Lebanese Studies at USEK.
  • A team of a historian and machine learning researchers at NC State.
  • An assistant professor and historian at Lebanese Association for History.
@@
E.2 Composition
@@
The dataset does not contain any confidential information. The public portion of the dataset includes page images that are already in the public domain. For the restricted part of the dataset, the images are proprietary and require permission from the owner for use only in non-commercial research.
@@ E.3 Collection process
E.4 Preprocessing/cleaning/labeling
@@
  • The original SFR code is available at: https://github.com/cwig/start_follow_read
  • The SFR code adapted for Arabic is available at https://github.com/mehreenmehreen/muharaf link.
E.5 Uses
@@ E.7 Maintenance
Table 4: Collections in the Muharaf dataset. Each image’s filename starts with the string shown in the “Prefix” column.

### v2

@@ 1 Introduction
The rest of this paper is organized as follows. Section 2 outlines the relevant characteristics of the Arabic script and Section 3 overviews the existing Arabic datasets for OCR/HTR. Section 4 describes the pipeline for collecting data and Section 5 details the dataset features, formats, and statistics. Section 6 provides the baseline results from our preliminary experiments on HTR. Section 7 concludes the paper and discusses the limitations of this work along with future directions.
2 Characteristics of Arabic script in Muharaf
The Arabic alphabet is believed to have its roots in the Nabatean alphabet from the third century [12]. The study of the origins of Arabic script and its evolution from classical Arabic to modern standard Arabic is a huge undertaking and still a subject of open scientific study and debate [33]. For example, the early Arabic writings did not contain dots (ijam) but later the script evolved to include them. The Muharaf dataset has writing samples from the early 19th century to the early 21st century, a period during which the use of dots was well established. Most of Muharaf’s samples are in the Arabic script Ruq‘ah. Ruq‘ah became the most common writing style in the late Ottoman period and early post-Ottoman period in the area we know today as the Middle East (which includes Egypt, but not North Africa).
A few challenges related to OCR/HTR of the Arabic script include the following:
  1. 1.
    The shape of each character of the Arabic script depends upon its contextual position within a word. Many characters have four different shapes depending upon whether they are in their isolated, initial (at the beginning of a word), medial (in the middle of a word), or final forms (at the end of a word). This poses a challenge for the OCR/HTR system that has to recognize all the different forms of the same character. Table 2 shows an example of 4 Arabic characters and their different forms.
  2. 2.
    Additional symbols in Arabic script include ijam, which are dots present above or below a character. Two characters may have the same basic shape but different numbers of dots to tell them apart. For example, the ijam or dots distinguish the bā’ (one dot below) from tā’ (two dots above), as shown in Table 2. An HTR system may misclassify a character because of the ijam.
  3. 3.
[…]
Table 1: Isolated, initial, medial, and final forms of 4 different letters from the Arabic alphabet.
[…]
[separator=pipe, tabular=C1.7cm C1.3cm C1.3cm C1.3cm C1.3cm, table head=Letter Isolated FormInitial Position Medial PositionFinal Position
, head to column names, late after last line=
[…]
@@ 3 Publicly available Arabic resources
The number of publicly available Arabic datasets is less than those available for languages written in the Latin script. Many Arabic datasets were tailored for specialized tasks, e.g., BADAM for baseline detection [14], HADARA80P for word spotting [26], AHDB for detecting and recognizing numbers on legal checks [4], and WAHD for writer identification [3]. In our subsequent discussion, we focus on offline, handwritten datasets for Arabic HTR, where ground truths for text are available. We may classify offline handwritten HTR datasets into two categories:
@@
[separator=pipe, tabular=C1.7cm C1.1cm C1.1cm L1cm L6.9cm, table head=Dataset Category Annotated Text Lines Total Writers Vocabulary/Composition
@@
a From unique subject IDs in scribe_demographic file.
b Based on text-independent features for handwriting analysis [27].
- Total number of writers is unspecified or unknown due to the nature of the dataset.
@@
Table 3: A comparison of various HTR datasets in terms of total pages, text regions, and total lines.
@@
Contrasting features of Muharaf and other available Arabic datasets
The existing Arabic datasets are a valuable resource for the research community. Muharaf supplements them with its own unique features and characteristics, discussed below:
  1. 1.
    As shown in Table 3, Muharaf has the largest number of line images as compared to IAM, RASAM, RASM, and KHATT.
  2. 2.
    KHATT and MADCAT datasets are category 1 (scribed) datasets, where writers were given text to write under controlled experimental conditions. Muharaf is a category 2 (original) dataset, where original historic manuscripts were scanned and transcribed.
  3. 3.
    RASAM and RASM are two Arabic datasets of scanned original historic manuscripts from category 2. The main difference between these two datasets and Muharaf lies in the Arabic script with which they were written (see points 4 and 5 below). Another difference is that RASAM and RASM handwritten pages belong to books, with calligraphic handwriting that is very neat and uniform across pages, written in straight horizontal lines. They were written by scholars in their respective fields. Muharaf includes informal/personal styles of writing, which were very common in the 19th and 20th centuries. The samples vary from very neat to barely legible writing. The handwriting samples of the same individual can be different over different documents or letters. Moreover, the text lines can be slanted upward or downward instead of horizontally straight lines.
  4. 4.
    RASAM has three types of manuscripts from the 10th century. They are scanned pages of books, which were written in the “Meghrebi script” also known as the “Round script” [33]. As the name suggests, this script has very rounded shaped letters. In contrast, Muharaf’s documents are mostly Ruq‘ah script, which is used for everyday or casual writing. It is composed of straight, short lines, and simple curves.
  5. 5.
    RASM has 4 different types of manuscripts of scientific writings from the 8th century to the 19th century. Muharaf has images from 50 different collections, each collection having one or more writers. Like RASAM, RASM’s handwriting styles are calligraphic, very neat, and uniform across all pages as opposed to Muharaf, where the writer may not have very careful or readable handwriting.
The OCR of text from different Arabic datasets like RASAM and RASM has its own challenges and is by no means an easy feat. We intend to supplement the existing Arabic datasets with a variety of handwritten images with the goal of digitizing handwritten documents from the late 19th century to the mid-20th century.
@@ 4 Data collection process
We developed an image-labeling software named ScribeArabic [^2] For code and a link to demo, please see https://github.com/MehreenMehreen/ScribeArabic specialized for annotating and transcribing Arabic page images of the Muharaf dataset. This software allows a user to annotate text lines in a browser window and transcribe them in a panel next to it. A separate module has the option to label various page elements. A screenshot of this software is shown in Figure 2. We have made the source code for ScribeArabic publicly available (see Section D of the supplementary material for all repository links).
@@
We assembled a team of expert Arabic speakers to transcribe historic Arabic manuscripts with technical support from machine learning researchers. The text lines of the first 180 images of our curated dataset were annotated by non-Arabic speakers using the ScribeArabic software. A Lebanese Arabic history professor then manually entered the transcriptions into an Excel sheet. We chose Excel for entering transcriptions because initially ScribeArabic supported only line annotations. For the next 1400140014001400+ images, we used our upgraded ScribeArabic software to allow the direct input of transcriptions in a browser window. The annotations and transcriptions for these 1400140014001400+ images were performed by two native Lebanese Arabic speakers adept at archival Arabic. Their transcriptions were checked by a third Lebanese Arabic expert who was also a historian. Section F.1 on page F.1 has more details of the qualifications of the annotation and transcription team.
@@ 5.1 PAGE-XML format for OCR/HTR datasets
The PAGE-XML format is an XML-based page image representation framework introduced by the PRImA Research Lab [31]. This format incorporates various image characteristics as well as the information on page layout and its contents at various levels of granularity. The official, full description of the XML schema is hosted at https://www.primaresearch.org. A conceptual class diagram of the subset of the PAGE-XML hierarchy used in the Muharaf dataset is shown in Figure 3. Where applicable, the following labeled regions are present in the PAGE-XML file of each image (see Section D.3 of the supplementary material for illustrative examples).
@@ 6 Baseline HTR results and discussion
Table 4: HTR results repeated over three random splits of Muharaf. Both page level and line level results are included.
@@ 6.1 Evolution of HTR performance as more data were collected
7 Conclusion and Future Work
@@
To the best of our knowledge, Muharaf is the largest publicly available Arabic dataset comprising fully annotated and transcribed historical manuscript pages at the text-line level. However, this collection is not devoid of limitations. The process of identifying all the writers and the exact timeline for each document will continue after the initial release of the dataset. For documents where the writer’s information cannot be extracted, e.g., the case where a scribe penned a legal document or a church record, one may want to define categories of different writers and writing styles. Another area of interest is the use of Muharaf transcriptions for the extraction of linguistic knowledge and the identification of the colloquial form of the Arabic language used in a particular period. Language models based on this information may improve the performance of the HTR system. We invite researchers in the document analysis and OCR community to utilize this dataset and advance the state of the art.
@@
We thank Stephen Randall Filios from Family Search for initiating discussions and providing feedback on tagging page elements in document images. We thank Elham Abdallah who is the Assistant University Librarian at USEK for providing support and coordinating work between NC State and USEK.
We acknowledge the computing resources provided by North Carolina State University High-Performance Computing Services Core Facility (RRID:SCR_022168). We also thank Andrew Petersen for his assistance and technical guidance on running jobs on the HPC.
@@ Appendix C Procedure for generating keywords for each collection
You will be given Arabic text. Based on the text, assign it the most three
relevant keywords from this list:
@@ 4. Give output as JSON format with keys:
Note that the prompt and the keywords present above were manually refined iteratively based on the output of GPT. We generated these keywords for the general interest of the research community who are non-Arabic speakers. However, we do not claim that these summaries and keywords are 100% correct.
@@ D.3 PAGE-XML converter and page elements viewer
We include the code for converting ScribeArabic’s JSON files to PAGE-XML files using a PAGE-XML converter. Instructions for downloading the code and running it are at https://github.com/mehreenmehreen/xml_converter. Users can also download a custom viewer for inspecting the annotated page elements on a document page. The viewer shows the different annotated page elements in different colors. A screenshot of this viewer is shown in Figure 6(a). This app was written in Python using the Tkinter library. Figure 6(b) illustrates examples of various page elements described in Section 5.1.
Figure 6: (a) Viewer of the annotated page elements showing floating regions, a paragraph region, and a signature region. Individual text lines are enclosed in black polygonal boundaries. (b) Various examples of different page elements including main paragraph regions, floating regions, a page number region, a signature region, and a graphics region.
@@ D.4 Start, Follow, Read — Arabic
  • Additional inference results from training the Muharaf dataset and its public portion. There are also additional cross-dataset evaluation results.
D.5 Pretraining the model
The SOL and LF networks of SFR-Arabic were trained from scratch with random weights. We pretrained the HW model using synthetically generated Arabic lines. The text of the Arabic lines was taken from the Khaleej-2004 newspaper corpus [1] and Watan-2004 corpus [2]. We have provided the details of the pretrained model on our Github webpage with a link to download it.
D.6 License
@@
D.7 Ethics statement: privacy, legal, or ethical issues
Muharaf-public has manuscript images from the archives of Phoenix Center for Lebanese Studies at USEK and KCLDS at NC State. These images were already publicly accessible before our proposed redistribution effort. Readers could access the images through the centers’ websites or as hard copies through the libraries. The research process described in this manuscript sticks to the ethical standards, and there are also no concerns regarding the leakage of personal information.
Muharaf-restricted is distributed under a restricted license. The limited licensing is due to the proprietary nature of the images, and hence are authorized to use only with permission from the owners. The usage restrictions include that the dataset shall be used solely for research purposes and not be redistributed without permission. We will require researchers to agree to the statement of ethical use of data as a requirement of downloading the restricted part of the dataset.
Appendix E Additional comparisons and cross-dataset validation results
Table 6 shows the comparison of SFR with a transformer based system TrOCR [17] on the IAM dataset [19]. While the performance of TrOCR is better than SFR, it has a higher computational complexity and requires more resources. Our choice of SFR is based on the goal of developing a system for a low-resource environment. Each network in this system can fit on an 8 GB card (mentioned in Section 6.1), making this system ideal for a low-resource environment. State-of-the-art models like transformers cannot be realistically deployed in a resource-constrained setting. By choosing a more traditional CNN-based network, we are willing to trade a slight reduction in accuracy for a smaller, less resource-intensive model.
[…]
Table 6: Comparison of TrOCR [17] and SFR [35] on the IAM dataset.
[…]
[separator=pipe, tabular= L4.5cm C2cm C1.5cm , table head=Model CER (%)Parameters
, head to column names, late after line=
[…]
Table 7: HTR performance of systems on KHATT dataset trained with different datasets. The test images are the line images from the test set in KHATT’s “unique paragraph” directory.
[…]
[separator=pipe, tabular= L5.5cm C3.5cm C1.2cm C1.3cm, table head=Model Training DatasetCER (%)Parameters
, head to column names, late after line=
[…]
Appendix F Datasheet for Muharaf Dataset
@@ F.1 Motivation
A majority of the annotations and transcriptions (1400+ images) were completed by Arabic speakers who manage digital archives of Arabic manuscripts and facilitate information access from a large number of Arabic documents. Their exact designations are:
  • Carlos Younes: Head of the reference and external relations at Phoenix Center for Lebanese-Historical Archives (PCLS), USEK.
  • Georges Habchi: Head of Storage Division at Digital Development Center, USEK.
The transcriptions of both individuals were checked by a full professor of history who is also a historian and the director of USEK Library, Phoenix Center for Lebanese Studies.
Some transcriptions (around 180 images) were completed by:
  • Amin Elias: Assistant professor teaching history at the center for Lebanese studies.
These transcriptions were checked by a full professor of history at KCLDS, NC State. Each image of the packaged Muharaf dataset is accompanied by a JSON and XML file. The JSON and XML files contain the names of the annotator and the transcriber of that image. Table 8 summarizes the total number of images transcribed by each team member.
@@
Table 8: Summary of total images transcribed by each member of the transcription team.
[…]
[separator=pipe, tabular= L5.5cm C3.5cm, table head=Transcriber Total images transcribed
, head to column names, late after line=
[…]
F.2 Composition
@@
The dataset does not contain any confidential information. The public portion of the dataset includes page images that are already publicly accessible. For the restricted part of the dataset, the images are proprietary and require permission from the owner for use only in non-commercial research.
@@ F.3 Collection process
Effectiveness of QA process
To assess the effectiveness of the QA process, we evaluated the CER and WER of several batches of received transcriptions both before and after the QA process. The results are summarized in Table 9. As shown, only minimal changes were made to the transcriptions during the QA phase.
@@
Table 9: CER and WER of transcriptions of 5 different batches before and after the QA phase.
[…]
[separator=pipe, tabular= L2.5cm L1.5cmL1.5cmC1.5cm, table head=Date Batch Received (2024) CER (%) WER (%)Total Images
, head to column names, late after line=
[…]
F.4 Preprocessing/cleaning/labeling
@@
F.5 Uses
@@ F.7 Maintenance
Table 5: Collections in the Muharaf dataset. Each image’s filename starts with the string shown in the “Prefix” column. The time period is listed if known.